1. graph the equation and identify the y-intercept. x + 5y = 20

2. The valve, v, of a shopkeeper's inventory software program is given by v = -1/2t + 6, where t is the number of year's since the shopkeeper bought the program. What is the program worth 4 years after it is first purchased?
Plot the point with the given t-valve, 4.
3. Find the y-intercept and x-intercept for the equation 3x + 2y = 12 What is the y-intercept? Type an ordered pair.
4. Find the intercepts and then use them to graph the equation 3y - 9 = 3x

Answer by Alan3354(69443)

1. graph the equation and identify the y-intercept. x + 5y = 20
At the y-intercept, x = 0
0 + 5y = 20
y = 4
------------------
2. The valve, v, of a shopkeeper's inventory software program is given by v = -1/2t + 6, where t is the number of year's since the shopkeeper bought the program. What is the program worth 4 years after it is first purchased?
Plot the point with the given t-valve, 4.
v(t) = -t/2 + 6
v(4) = 4
-----------------------
3. Find the y-intercept and x-intercept for the equation 3x + 2y = 12 What is the y-intercept? Type an ordered pair.
Do it like #1.
-------------------
4. Find the intercepts and then use them to graph the equation 3y - 9 = 3x
Similar to #1.
-------------------
To graph linear eqns, you need 2 points. Pick values for x or y and find the 2 points. Plot them, and draw a line thru them.